Gallium arsenide has a zinc blende crystal structure. Assume a lattice constant

of a = 5.65 Å. Determine the density of gallium arsenide in g/cm3.


1
Expert's answer
2021-07-09T05:15:24-0400

8 atoms per unit cell: 4 Ga and 4 As

atomic weights:

Ga = 69.72 Da or u

As = 74.92

1u = 1.660 ×10−27 kg

Total mass in unit cell: M = 4 × (69.72 + 74.92) ×1.660 ×10−27 kg M = 9.60 ×10−25 kg

Density = mass/volume:

ρ = 9.60 ×10-25 kg/

(5.65×10-10 )3

= 5.32 ×103 kg/m3

(These are proper MKS units. Also called International System of Units (SI).)

ρ = 5.32 ×103 kg/

m3 ×

103

gm/

kg

×

1/

102

( cm/ m)3

= 5.32 gm/cm3

ρ = 5.32 gm/cm3